OpenStudy (anonymous):

A 0.4791g sample was analyzed and was found to contain the following masses of elements: Carbon, 0.1929g; Hydrogen, 0.01079g; Oxygen, 0.08566g; Chlorine, 0.1898g. Determine the empirical formula of the compound. This is the first question of the assignment, and if anyone could help me solve this and explain the process to me that would be helpful too

OpenStudy (rogue):

I don't have the time to write a full solution, but here are the basic steps you need to follow: 1. Divide the masses of each of the elements in the sample by their molar masses to get the moles of each elements in the sample. 2. Compare the ratio of each element to one another to find the formula of the molecule.

OpenStudy (rogue):

Here is a better guide with examples: http://dl.clackamas.cc.or.us/ch104-04/empirica.htm

OpenStudy (rogue):

ratio of moles*

OpenStudy (xishem):

To give support to Rogue, here is the work done:\[0.1929g(C)*\frac{1mol(C)}{12.01g(C)}=0.01606mol(C)\]\[0.01079g(H)*\frac{1mol(H)}{1.008g(H)}=0.01070mol(H)\]\[0.08566g(O)*\frac{1mol(O)}{16.00g(O)}=0.005354mol(O)\]\[0.1898g(Cl)*\frac{1mol(Cl)}{35.45g(Cl)}=0.005354mol(Cl)\]Now express each element as a ratio to an element with the fewest number of moles, and round to the nearest whole number: \[mols(C)=\frac{0.01606}{0.005354}=3\]\[mols(H)=\frac{0.01070}{0.005354}=2\]\[mols(O)=\frac{0.005354}{0.005354}=1\]\[mols(Cl)=\frac{0.005354}{0.005354}=1\]Based on that ratio of moles, what do you think the final empirical formula is?

OpenStudy (anonymous):

Thank you both for the answers :-) both extremely helpful